OpenStudy (anonymous):

how do you solve 8x-2y+=-10

OpenStudy (idealist10):

For which variable?

OpenStudy (anonymous):

8x+10=2y y=4x+5

OpenStudy (anonymous):

or 8x=2y-10 x=1/4x-5/4

OpenStudy (anonymous):

Thanks everyone